A bus travels a certain distance with a speed 30 m/sec and returns with a speed 20 m/sec then find average speed during total journey. 

Grade:7

2 Answers

Vikas TU
7 years ago
Let t1 be the time to travel x distance with 30 m/sec.
Therfore, 
t1 = x/30 …................(1)
 
 and t2 be the time to travel x distance with 20 m/sec.
t2 = x/20 ….........(2)
 
Average speed = Total Distance/Total time
= 2x/(t1 + t2)
=2x(x/30 + x/20) {from eqns. (1)  and (2))
= 24 m/s {x gets cancelled out}
 
Ajay Choudhury
7 years ago
let the distance be x
thus time taken to go=x/30
time taken to come=x/20
average speed=120x/5x
                     =24m/s